OpenStudy (anonymous):

Simplify: (4x − 6) + (3x + 6) 7x 7x − 12 x 7x + 12

OpenStudy (anonymous):

7x

OpenStudy (foolaroundmath):

(4x - 6) + (3x + 6) = 4x+ 3x - 6 + 6 = x(4+3) + 6-6 = 7x

OpenStudy (anonymous):

thank you :)